OpenStudy (anonymous):

What are the sine, cosine, and tangent of Θ = 7 pi over 4 radians?

OpenStudy (anonymous):

\[\sin \frac{ 7 \pi }{ 4 }=\sin \left( \pi-\frac{ \pi }{ 4 } \right)=\sin \frac{ \pi }{ 4 }=\frac{ 1 }{ \sqrt{2} }\] \[\cos \frac{ 7 \pi }{ 4 }=\cos \left( \pi-\frac{ \pi }{ 4 } \right)=-\cos \frac{ \pi }{ 4 }=-\frac{ 1 }{ \sqrt{2} }\] \[\tan \frac{ 7 \pi }{ 4 }=\tan \left( \pi-\frac{ \pi }{ 4 } \right)=?\]